# A variational approach to dimension-free self-normalized concentration

### v1

@@ 2 Preliminaries
Let $\Theta$ be a measurable parameter space. For each $\theta\in\Theta$, let $Z(\theta)\equiv(Z_{t}(\theta))_{t\geq 0}$ be a stochastic process upper bounded by a nonnegative supermartingale $Y(\theta)\equiv(Y_{t}(\theta))_{t\geq 0}$. Assume that all processes are adapted to the same filtration $\mathcal{F}$ and that $Y_{0}(\theta)\leq 1$ for each $\theta\in\Theta$. Let $\nu$ be a data-free distribution over $\Theta$. Then, for all $\delta\in(0,1)$, with probability $1-\delta$, for any $\mathcal{F}$-adapted stopping time $\tau$ and any $\mathcal{F}_{\tau}$-measurable distribution $\rho_{\tau}$ over $\Theta$,
@@
Notation.
We denote the eigenvalues of a matrix $M$ as $\gamma_{\min}(M)=\gamma_{1}(M)\leq\dots\leq\gamma_{d}(M)=\gamma_{\max}(M)$. We let $\|M\|_{{\textnormal{op}}}=\gamma_{\max}(M)$ be the operator norm of $M$. For a convex function $f:I\to\mathbb{R}$, $I\subseteq\mathbb{R}$, we denote its convex conjugate by $f^{*}(y)=\sup_{x\in I}(xy-f(x))$, which is also a convex function. The convex conjugate is also called the Legendre–Fenchel transform. We let $\mathbb{N}_{0}=\{0,1,2,\dots\}$ denote the natural numbers including zero and $\mathbb{R}_{>0}$ denote the positive real numbers. We use $A\succeq B$ to signify that $A$ is greater than $B$ in the Loewner order, i.e., ABA-B is positive semidefinite. For a vector $v$ and positive semidefinite matrix $A$, $\|v\|_{A}$ denotes the norm defined as $\|v\|_{A}^{2}=\langle v,Av\rangle$, where $\langle\cdot,\cdot\rangle$ is the usual dot product in $\mathbb{R}^{d}$ unless otherwise specified. We let $a\vee b=\max\{a,b\}$ and use $A^{\intercal}$ to refer to the transpose of the matrix $A$. Throughout, the indices on the sum $\sum_{j\leq t}$ should be assumed to run from 1 to $t$.
@@ 3 Warmup: Sub-Gaussian Processes
Let us begin by considering the special case of sub-Gaussian processes [^7] It’s perhaps worth mentioning “sub-Gaussian process” has a double meaning in the literature. In this paper, a sub-Gaussian process refers to Definition 2.1 with $\psi=\psi_{N}$. This is different than the sub-Gaussian process studied by Talagrand [55], which is a collection of zero-mean random variable $\{Y_{\theta}:\theta\in\Theta\}$ such that $\mathbb{E}[\exp(\lambda(X_{\theta}-X_{\phi}))]\leq\exp(\psi_{N}(\lambda)\rho^{2}(\theta,\phi))$ for some metric $\rho$ and all $\theta,\phi\in\Theta$. . That is, we assume that $\psi(\lambda)=\psi_{N}(\lambda)=\lambda^{2}/2$. This allows for a particularly clean application of the variational template, since we may take $\Theta=\mathbb{R}^{d}$ as our parameter space. (For $\psi=\psi_{N}$, one can replace $\mathbb{S}^{d-1}$ with $\mathbb{R}^{d}$.) This in turn enables the use of normal distributions in Proposition 2.2.
@@
Theorem 3.1 recovers the bound of Abbasi-Yadkori et al. [1] if $(X_{t})_{t\geq 1}\subseteq\mathbb{R}^{d}$ is an $\mathcal{F}$-predictable process ($X_{t}$ is the $t$-th action taken by a bandit algorithm in their case) and $S_{t}=\sum_{k=1}^{t}\eta_{k}X_{k}$ where $\eta_{k}$ is scalar $\sigma$-sub-Gaussian noise. [^8] That is, $\mathbb{E}[\exp(\lambda\eta_{k})|\mathcal{F}_{k-1}]\leq\exp(\lambda^{2}\sigma^{2}/2)$ Then $(S_{t})$ is sub-$\psi_{N}$ with variance proxy $(V_{t})$ defined as $V_{t}=\sigma^{2}\sum_{k=1}^{t}X_{k}X_{k}^{\intercal}$. Taking $U_{0}=\sigma^{2}\rho V_{0}$ for some $V_{0}$ and rearranging (3.1) gives Theorem 1 of Abbasi-Yadkori et al. [1]. More generally, Theorem 3.1 applies to any sub-Gaussian process described in Section 2, such as $S_{t}=\sum_{k\leq t}X_{k}$ and $V_{t}=\sum_{k\leq t}X_{k}X_{k}^{\intercal}$ when $X_{t}$ is conditionally symmetric.
@@ 4 General Sub-$\psi$ Processes
Theorem 3.1 relied on transforming the sub-$\psi$ condition from one which holds over $\mathbb{S}^{d-1}$ to one which holds over $\mathbb{R}^{d}$, thus enabling us to take $\Theta=\mathbb{R}^{d}$ and use normal distributions in Proposition 2.2. This trick cannot be performed for general sub-$\psi$ processes because $\lambda_{\max}$ may be finite. However, for general $\psi$ functions we may replace $\mathbb{S}^{d-1}$ with $\mathbb{B}^{d}$, the unit ball in $\mathbb{R}^{d}$. [^9] More specifically, we can upper bound any $\psi$ with some $\widehat{\psi}$ that allows us to replace $\mathbb{S}^{d-1}$ with $\mathbb{B}^{d}$. Such $\widehat{\psi}$ are called super-Gaussian. See Remark 4.2 for further discussion. This allows us to use nested uniform distributions over $\Theta=\mathbb{B}^{d}$ in Proposition 2.2.
The machinery we use to handle general sub-$\psi$ processes does have drawbacks over what was used in Section 3 to study the specific case of sub-Gaussian processes. For one, when we apply the results in this section to sub-Gaussian processes, the resulting bound is looser than Theorem 3.1 (see Figure 1). Second, the parameter $\lambda$ of the sub-Gaussian process in Theorem 3.1 can be optimized independently of other terms in the bound. Such closed-form optimization of $\lambda$ is not possible in the more general case. We instead end up with line-crossing inequalities: For any given $\lambda$ we obtain a bound on the probability that $\|S_{t}\|_{V_{t}^{-1}}$ ever crosses a threshold parameterized by $\lambda$. Different values of $\lambda$ result in a bound which is tighter at different values of $\det V_{t}$. Theorem 4.1 gives the line-crossing inequality for general sub-$\psi$ processes.
Line-crossing inequalities are common in works studying the concentration of sub-$\psi$ processes [33, 34]. To go from line-crossing inequalities to bounds which are tight at all (intrinsic) times, we deploy a technique which has come to be known as “stitching,” though see Section 4.1 for more background. We provide two kinds of stitched bounds: One for general sub-$\psi$ processes, and one specifically tailored for sub-Gamma processes.
We must define two quantities before stating our first result. First, for a positive semidefinite matrices $M_{1}$ and $M_{2}$, define their Rayleigh-Ritz maximum [^10] This is also the maximum over the generalized Rayleigh coefficient between $M_{1}$ and $M_{2}$, and the solution to the generalized eigenvalue problem $M_{1}x=\lambda M_{2}x$.
@@
When using the Rayleigh-Ritz maximum hereafter, we will always consider matrices $M_{1},M_{2}$ such that $M_{2}\succeq M_{1}$, thus implying that $0\leq\alpha(M_{1},M_{2})\leq 1$. Next, given a sub-$\psi$ process $(S_{t},V_{t})$ and a matrix $U_{0}$ such that $V_{t}\succeq U_{0}$ for all $t\geq 1$, let
$\alpha_{t}=\alpha(U_{0},V_{t}),$ (4.2)
and define the sequence of functions $(g_{\psi,t}:\textnormal{Im}(\psi)\to[0,1])_{t\geq 1}$ by
$g_{\psi,t}(z)=\sqrt{\alpha_{t}\psi^{-1}(z)+1-\alpha_{t}}-\sqrt{\alpha_{t}\psi^{-1}(z)}.$ (4.3)
These functions may appear unwieldy at first glance but they behave relatively simply. Indeed, as $\alpha_{t}\to 0$ (which is the case if $\gamma_{\min}(V_{t})\to\infty$), $g_{\psi,t}(z)\to 1$ with a rate dictated by $\psi^{-1}$. For most common sub-$\psi$ processes and matrices $V_{t}$ which have at least some spread in all directions, $g_{\psi,t}(z)$ is above 0.5 after 100 samples for all $\lambda\in\textnormal{Im}(\psi)$. See Figure 1.
We now present the line-crossing inequality for sub-$\psi$ processes. The proof is in Appendix B.2.
@@
Let $(S_{t},V_{t})$ be a sub-$\psi$ process such that $\lambda\mapsto\psi(\lambda)/\lambda^{2}$ is nondecreasing. Let $U_{0}$ be a positive definite matrix such that $V_{t}\succeq U_{0}$ for all $t\geq 1$. Let $\alpha_{t}=\alpha_{t}(U_{0},V_{t})$ be the Rayleigh-Ritz maximum between $U_{0}$ and $V_{t}$. Then, for any $\delta\in(0,1)$ and any $\lambda\in[\psi^{-1}(1/\gamma_{\min}(U_{0})),\lambda_{\max})\cap\textnormal{Im}(\psi)$, with probability $1-\delta$, for any stopping time $\tau$,
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\frac{\|U_{0}\|_{\textnormal{op}}^{1/2}}{\lambda g_{\psi,\tau}(\lambda)}\left(\frac{1}{2}\log\left(\frac{\det V_{\tau}}{\det U_{0}}\right)+1+\log(1/\delta)\right)+\frac{g_{\psi,\tau}(\lambda)\psi(\lambda)}{\lambda\|U_{0}\|_{\textnormal{op}}^{1/2}}.$ (4.4)
@@
Theorem 4.1 assumes that $\psi(\lambda)/\lambda^{2}$ is a nondecreasing function, a condition that is sometimes referred to as $\psi$ being super-Gaussian. This assumption is without loss of generality in the following sense: As we detailed in Section 2, any $\psi$ function can be upper bounded by a constant times $\psi_{G,c}$, which is super-Gaussian. (In fact, all sub-$\psi$ processes mentioned in Section 2 are super-Gaussian). Thus, Theorem 4.1 may be applied to any sub-$\psi$ process, at the cost of some looseness in constants if $\psi$ is strictly sub-Gaussian (e.g., sub-Bernoulli).
@@
Figure 1: Left: The growth of $g_{t}$ for various $\psi$ functions and across various values of $\lambda$. For sub-gamma, sub-Poisson, and sub-exponential we fix $c=1/4$. Right: A comparison of Theorem 3.1 and Theorem 4.1 instantiated for $\psi=\psi_{N}$. In both figures we use $U_{0}=I_{d}$ and $V_{t}=U_{0}+\sum_{k\leq t}X_{s}X_{s}^{\intercal}$ where the vectors $X_{t}$ are chosen based on a bandit algorithm. Details may be found in Appendix C.
@@
Several remarks are in order. First, the requirement that $\lambda\geq\psi^{-1}(1/\gamma_{\min}(U_{0}))$ comes from our proof technique. More specifically, we use a prior which is uniform over the ellipsoid $\{\theta:\psi(\lambda)\theta^{t}U_{0}\theta\leq 1\}$, which we require to be a subset of $\mathbb{B}^{d}$. This holds if $\psi(\lambda)U_{0}\succeq I_{d}$, i.e., $\lambda\geq\psi^{-1}(\gamma_{\min}^{-1}(U_{0}))$. The requirement that $\lambda\leq\lambda_{\max}$ comes from Definition 2.1, and the requirement that $\lambda\in\textnormal{Im}(\psi)$ comes from the definition of $g_{\psi,t}$. Note, however, that for common $\psi$ functions including $\psi_{E}$, $\psi_{P}$, $\psi_{G}$ and $\psi_{N}$, $\textnormal{Im}(\psi)=\mathbb{R}_{\geq 0}$, so requiring that $\lambda\in\textnormal{Im}(\psi)$ is often a vacuous constraint.
@@
Second, Theorem 4.1 may appear as though it has the wrong order of magnitude. Indeed, Theorem 3.1 is a bound on $\|S_{\tau}\|_{V_{\tau}^{-1}}^{2}$, whereas (4.4) is a bound on $\|S_{\tau}\|_{V_{\tau}^{-1}}$ (without the square). To set yourself at ease, let $D_{t}(\delta)=\frac{1}{2}\log(\det V_{t}/\det U_{0})+1+\log(1/\delta)$ and assume for simplicity that $\|U_{0}\|=1$. Then, in the sub-Gaussian case, (4.4) reads
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\frac{D_{\tau}(\delta)}{\lambda g_{\tau}(\lambda)}+\frac{\lambda g_{\tau}(\lambda)}{2}.$
@@
For a fixed $t$, if $\lambda$ is such that $\lambda g_{t}(\lambda)=\sqrt{2D_{t}(\delta)}$ (note that the map $x\mapsto xg_{\psi_{N},t}(x)$ is surjective on $\mathbb{R}_{\geq 0}$), in which case we obtain $\|S_{t}\|_{V_{t}^{-1}}^{2}\leq D_{t}(\delta)$. This matches Theorem 3.1 up to an additive factor of 2. However, this only demonstrates that there exists a $\lambda$ which makes the widths roughly equal. We cannot in fact choose $\lambda$ such that $\lambda g_{t}(\lambda)=\sqrt{2D_{t}(\delta)}$ since $\lambda$ cannot be data-dependent. Moreover, such a choice of $\lambda$ would only ensure the widths are comparable at a particular fixed time $t$.
@@
If $\lambda$ is not optimized, how does Theorem 4.1 with $\psi=\psi_{N}$ compare to Theorem 3.1? The right hand panel of Figure 1 plots the comparison for various values of $\lambda$. As one would expect, Theorem 4.1 is looser since it handles a wider class of distributions and is not explicitly optimized for sub-Gaussian processes. That said, Theorem 4.1 loses surprisingly little over Theorem 3.1, even for fixed values of $\lambda$. As we can see, one faces a tradeoff when choosing $\lambda$, with smaller values being tighter for larger $t$ (though sometimes this tradeoff is minimal compared to the size of the bound; see Figure 2 for instance). Next we explore how to optimize $\lambda$ over time.

### v2

@@ 2 Preliminaries
Let $\Theta$ be a measurable parameter space. For each $\theta\in\Theta$, let $Z(\theta)\equiv(Z_{t}(\theta))_{t\geq 0}$ be a stochastic process upper bounded by a nonnegative supermartingale $L(\theta)\equiv(L_{t}(\theta))_{t\geq 0}$. Assume that all processes are adapted to the same filtration $\mathcal{F}$ and that $Y_{0}(\theta)\leq 1$ for each $\theta\in\Theta$. Let $\nu$ be a data-free distribution over $\Theta$. Then, for all $\delta\in(0,1)$, with probability $1-\delta$, for any $\mathcal{F}$-adapted stopping time $\tau$ and any $\mathcal{F}_{\tau}$-measurable distribution $\rho_{\tau}$ over $\Theta$,
@@
Notation. We denote the eigenvalues of a matrix $M$ as $\gamma_{\min}(M)=\gamma_{1}(M)\leq\dots\leq\gamma_{d}(M)=\gamma_{\max}(M)$. We let $\|M\|_{{\textnormal{op}}}=\gamma_{\max}(M)$ be the operator norm of $M$. For a convex function $f:I\to\mathbb{R}$, $I\subseteq\mathbb{R}$, we denote its convex conjugate (Legendre–Fenchel transform) by $f^{*}(y)=\sup_{x\in I}(xy-f(x))$, which is also a convex function. We let $\mathbb{N}_{0}=\{0,1,2,\dots\}$ denote the natural numbers including zero and $\mathbb{R}_{>0}$ denote the positive real numbers. We use $A\succeq B$ to signify that $A$ is greater than $B$ in the Loewner order, i.e., ABA-B is positive semidefinite. For a vector $v$ and positive semidefinite matrix $A$, $\|v\|_{A}$ denotes the norm defined as $\|v\|_{A}^{2}=\langle v,Av\rangle$, where $\langle\cdot,\cdot\rangle$ is the usual dot product in $\mathbb{R}^{d}$ unless otherwise specified. We let $a\vee b=\max\{a,b\}$ and use $A^{\intercal}$ to refer to the transpose of the matrix $A$. Throughout, the indices on the sum $\sum_{j\leq t}$ should be assumed to run from 1 to $t$.
@@ 3 Warmup: Sub-Gaussian Processes
Let us begin by considering the special case of sub-Gaussian processes [^3] It’s perhaps worth mentioning “sub-Gaussian process” has a double meaning in the literature. In this paper, a sub-Gaussian process refers to Definition 2.1 with $\psi=\psi_{N}$. This is different than the sub-Gaussian process studied by Talagrand [57], which is a collection of zero-mean random variable $\{Y_{\theta}:\theta\in\Theta\}$ such that $\mathbb{E}[\exp(\lambda(X_{\theta}-X_{\phi}))]\leq\exp(\psi_{N}(\lambda)d^{2}(\theta,\phi))$ for some metric $d$ and all $\theta,\phi\in\Theta$. . That is, we assume that $\psi(\lambda)=\psi_{N}(\lambda)=\lambda^{2}/2$. This allows for a particularly clean application of the variational template, since we may take $\Theta=\mathbb{R}^{d}$ as our parameter space. (For $\psi=\psi_{N}$, one can replace $\mathbb{S}^{d-1}$ with $\mathbb{R}^{d}$.) This in turn enables the use of normal distributions in Proposition 2.2.
@@
To give a sense of how Proposition 2.2 is used to prove Theorem 3.1, let us give a brief outline of the proof here.
If $(S_{t},V_{t})$ is a sub-Gaussian process, then it follows that the process $(N_{t}(\theta))$ with $N_{t}(\theta)=\exp\{\lambda\langle\theta,S_{t}\rangle-\psi_{N}(\lambda)\langle\theta,V_{t}\theta\rangle\}$ is upper bounded by a supermartingale for all $\theta\in\mathbb{R}^{d}$ (not only for every $\theta\in\mathbb{S}^{d-1}$; this is because $\lambda_{\max}=\infty$). We may therefore apply Proposition 2.2 with $\Theta=\mathbb{R}^{d}$. In particular, we take $\rho_{\tau}$ to be Gaussian with mean $(V_{\tau}+U_{0})^{-1}S_{\tau}$ and covariance $(2\psi_{N}(\lambda))^{-1}(V_{\tau}+U_{0})^{-1}$, and $\nu$ to be Gaussian with mean 0 and covariance $(2\psi_{N}(\lambda))^{-1}U_{0}^{-1}$. (Recall that $\rho$ can be data-dependent while $\nu$ must be data-free.) Then $\int\log N_{t}(\theta)\textnormal{d}\rho_{\tau}=(\lambda-\psi_{N}(\lambda))\|S_{\tau}\|^{2}_{(V_{\tau}+U_{0})^{-1}}+\text{(extra terms)}$ and $D_{\textsc{kl}}(\rho_{\tau}\|\nu)=\frac{1}{2}\log(\det(V_{\tau}+U_{0})/\det U_{0})+\text{(the same extra terms)}$. Rearranging and taking $\lambda=1$ gives the result.
Arithmetic aside, the crux of the argument is noticing that we may work with $\Theta=\mathbb{R}^{d}$ instead of $\Theta=\mathbb{S}^{d-1}$, enabling us to choose $\rho$ and $\nu$ as Gaussians. This strategy unfortunately won’t be available to us for more general sub-$\psi$ distributions, which makes those settings more challenging. We will end up using uniform distributions instead.
Theorem 3.1 recovers the bound of Abbasi-Yadkori et al. [1] if $(X_{t})_{t\geq 1}\subseteq\mathbb{R}^{d}$ is an $\mathcal{F}$-predictable process ($X_{t}$ is the $t$-th action taken by a bandit algorithm in their case) and $S_{t}=\sum_{k=1}^{t}\eta_{k}X_{k}$ where $\eta_{k}$ is scalar $\sigma$-sub-Gaussian noise. Then $(S_{t})$ is sub-$\psi_{N}$ with variance proxy $(V_{t})$ defined as $V_{t}=\sigma^{2}\sum_{k=1}^{t}X_{k}X_{k}^{\intercal}$. Taking $U_{0}=\sigma^{2}V_{0}$ for some $V_{0}$ and rearranging (3.1) gives Theorem 1 of Abbasi-Yadkori et al. [1]. More generally, Theorem 3.1 applies to any sub-Gaussian process described in Section 2, such as $S_{t}=\sum_{k\leq t}X_{k}$ and $V_{t}=\sum_{k\leq t}X_{k}X_{k}^{\intercal}$ when $X_{t}$ is conditionally symmetric.
@@ 4 General Sub-$\psi$ Processes
Theorem 3.1 relied on transforming the sub-$\psi$ condition from one which holds over $\mathbb{S}^{d-1}$ to one which holds over $\mathbb{R}^{d}$, thus enabling us to take $\Theta=\mathbb{R}^{d}$ and use normal distributions in Proposition 2.2. This trick cannot be performed for general sub-$\psi$ processes because (in their definition) $\lambda_{\max}$ may be finite. However, for general $\psi$ functions we may replace $\mathbb{S}^{d-1}$ with $\mathbb{B}^{d}$, the unit ball in $\mathbb{R}^{d}$. More specifically, we can upper bound any CGF-like $\psi$ with some super-Gaussian $\widehat{\psi}$ that allows us to replace $\mathbb{S}^{d-1}$ with $\mathbb{B}^{d}$; see Remark 4.2 for further discussion. This allows us to use nested uniform distributions over $\Theta=\mathbb{B}^{d}$ in Proposition 2.2.
The machinery we use to handle general sub-$\psi$ processes does have drawbacks over what was used in Section 3 to study the specific case of sub-Gaussian processes. For one, when we apply the results in this section to sub-Gaussian processes, the resulting bound is looser than Theorem 3.1 (see Figure 1). Second, the parameter $\lambda$ of the sub-Gaussian process in Theorem 3.1 can be optimized independently of other terms in the bound. Such closed-form optimization of $\lambda$ is not possible in the more general case. We instead end up with line-crossing inequalities: for any given $\lambda$, we obtain a bound on the probability that $\|S_{t}\|_{V_{t}^{-1}}$ ever crosses a threshold parameterized by $\lambda$. Different (predetermined) values of $\lambda$ result in a bound which is tighter at different (predetermined) values of $\det V_{t}$. Theorem 4.1 gives the line-crossing inequality for general sub-$\psi$ processes.
Line-crossing inequalities are common in works studying the concentration of sub-$\psi$ processes [34, 35]. To go from line-crossing inequalities to bounds which are tight at all (intrinsic) times, we deploy a technique which has come to be known as “stitching” (see Section 4.2 for more background). We provide two kinds of stitched bounds: One for general sub-$\psi$ processes, and one specifically tailored for sub-Gamma processes. This will be the goal of Section 4.2.
4.1 A line-crossing inequality for sub-$\psi$ processes
We must define two quantities before stating our first result. First, for nonzero positive definite matrices $M_{1}$ and $M_{2}$, define their Rayleigh-Ritz quotient [^4] This is also the maximum over the generalized Rayleigh coefficient between $M_{1}$ and $M_{2}$, and the maximum solution to the generalized eigenvalue problem $M_{1}x=\lambda M_{2}x$.
@@ 4.1 A line-crossing inequality for sub-$\psi$ processes
When using the Rayleigh-Ritz quotient hereafter, we will always consider matrices $M_{1},M_{2}$ such that $M_{2}\succeq M_{1}$, thus implying that
$0\leq\alpha(M_{1},M_{2})\leq 1.$
Next, given a sub-$\psi$ process $(S_{t},V_{t})$ and a PSD matrix $U_{0}$, let
$\alpha_{t}=\alpha(U_{0},U_{0}+V_{t}),$ (4.2)
and for $t\geq 1$ define the function $g_{\psi,t}:\textnormal{Im}(\psi)\to[0,1]$ by
$g_{\psi,t}(z)=\sqrt{\psi^{-1}(z)}-\sqrt{\alpha_{t}\psi^{-1}(z)}.$ (4.3)
Clearly, $g_{\psi,t}(z)\to\sqrt{\psi^{-1}(z)}$ as $\alpha_{t}\to 0$, which occurs for example if $\gamma_{\min}(V_{t})\to\infty$. For most common sub-$\psi$ processes and matrices $V_{t}$ which have at least some spread in all directions, $g_{\psi,t}(z)$ is rapidly approaching $\sqrt{\psi^{-1}(z)}$ after 100 samples for all $\lambda\in\textnormal{Im}(\psi)$; see Figure 1. We now present the line-crossing inequality for sub-$\psi$ processes. The proof is in Appendix B.2.
@@
Let $(S_{t},V_{t})$ be a sub-$\psi$ process such that $\lambda\mapsto\psi(\lambda)/\lambda^{2}$ is nondecreasing. Let $U_{0}$ be positive definite and define the Rayleigh-Ritz quotient $\alpha_{t}=\alpha(U_{0},U_{0}+V_{t})$. Then, for any $\delta\in(0,1)$ and any $\lambda\in[\psi^{-1}(1/\gamma_{\min}(U_{0})),\lambda_{\max})\cap\textnormal{Im}(\psi)$, with probability $1-\delta$, for any stopping time $\tau$,
$\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}\leq\frac{1}{\lambda g_{\psi,\tau}(\lambda)}\left(\frac{1}{2}\log\left(\frac{\det(V_{\tau}+U_{0})}{\det U_{0}}\right)+1+\log(1/\delta)\right)+\frac{g_{\psi,\tau}(\lambda)\psi(\lambda)}{\lambda}.$ (4.4)
@@
Theorem 4.1 assumes that $\psi$ is super-Gaussian, i.e., that $\psi(\lambda)/\lambda^{2}$ is nondecreasing. This assumption is without loss of generality in the following sense: As we detailed in Section 2, any $\psi$ function can be upper bounded by a constant times $\psi_{G,c}$, which is super-Gaussian. (In fact, all sub-$\psi$ processes mentioned in Section 2 are super-Gaussian). Thus, Theorem 4.1 may be applied to any sub-$\psi$ process, at the cost of some looseness in constants if $\psi$ is strictly sub-Gaussian (e.g., sub-Bernoulli).
Let us provide a sketch of the proof of Theorem 4.1. Unlike the sub-Gaussian case (Theorem 3.1), we cannot convert the sub-$\psi$ condition into one that holds for all $\theta\in\mathbb{R}^{d}$. Due to the super-Gaussian assumption, however, we can assume that it holds for all $\theta\in\mathbb{B}^{d}$—the unit ball as opposed to the unit sphere.
We thus apply Proposition 2.2 to the process $Z_{t}(\theta)=\exp\{\lambda\langle\theta,S_{t}\rangle-\psi(\lambda)\langle\theta,(V_{t}+U_{0})\theta\rangle\}$ for $\theta\in\mathbb{B}^{d}$, where we take distributions $\nu$ and $\rho$ to be uniform over ellipsoids $\mathcal{E}_{\nu}$ and $\mathcal{E}_{\rho}$ sitting inside $\mathbb{B}^{d}$. The shape (hence volume) of $\mathcal{E}_{\nu}$ is a function of $U_{0}^{-1}$; that of $\mathcal{E}_{\rho}$ a function of $(V_{\tau}+U_{0})^{-1}$ (recall that $\rho$ can be $\mathcal{F}_{\tau}$-measurable). With these choices, integrating out $\int\log Z_{\tau}(\theta)\textnormal{d}\rho$ and applying Proposition 2.2 gives that with probability $1-\delta$,
$(\lambda\beta-\beta^{2}\psi(\lambda))\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}\leq D_{\textsc{kl}}(\rho\|\nu)+1+\log(1/\delta).$
The KL-divergence between $\rho$ and $\nu$ is the ratio of their volumes, which is $\frac{1}{2}\log(\det(V_{\tau}+U_{0})/\det U_{0})$, giving rise to the determinant-based bound. The parameter $\beta$ is a multipler on the mean of $\mathcal{E}_{\rho}$, ensuring that it sits inside $\mathcal{E}_{\nu}$. Optimizing $\beta$ gives the final result.
@@
Figure 1: Left: The growth of $g_{t}$ for various $\psi$ functions and across various values of $\lambda$. For sub-gamma, sub-Poisson, and sub-exponential we fix $c=1/4$. Right: A comparison of Theorem 3.1 (dotted black line) and Theorem 4.1 (various colored lines) instantiated for $\psi=\psi_{N}$. In both figures we use $U_{0}=I_{d}$ and $V_{t}=\sum_{k\leq t}X_{s}X_{s}^{\intercal}$ where the vectors $X_{t}$ are chosen based on a bandit algorithm. Details may be found in Appendix C.
@@
Several remarks on Theorem 4.1 are in order. First, the lower bound on $\lambda$, i.e., $\lambda\geq\psi^{-1}(1/\gamma_{\min}(U_{0}))$ comes from the proof technique described above. The precise form of $\mathcal{E}_{\nu}$ is $\{\theta:\psi(\lambda)\theta^{t}U_{0}\theta\leq 1\}$, which we require to be a subset of $\mathbb{B}^{d}$. This holds if $\psi(\lambda)U_{0}\succeq I_{d}$, i.e., $\lambda\geq\psi^{-1}(\gamma_{\min}^{-1}(U_{0}))$. Further, the requirement that $\lambda\leq\lambda_{\max}$ comes from Definition 2.1, and the requirement that $\lambda\in\textnormal{Im}(\psi)$ comes from the definition of $g_{\psi,t}$. Note, however, that for common $\psi$ functions including $\psi_{E}$, $\psi_{P}$, $\psi_{G}$ and $\psi_{N}$, $\textnormal{Im}(\psi)=\mathbb{R}_{\geq 0}$, so requiring that $\lambda\in\textnormal{Im}(\psi)$ is often a vacuous constraint.
@@
Second, Theorem 4.1 may appear as though it has the wrong order of magnitude. Indeed, Theorem 3.1 is a bound on $\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}^{2}$, whereas (4.4) is a bound on $\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}$ (without the square). To set ourselves at ease, let $D_{t}(\delta)=\frac{1}{2}\log(\det(V_{t}+U_{0})/\det U_{0})+1+\log(1/\delta)$. The bound in (4.4) reads
$\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}\leq\frac{D_{\tau}(\delta)}{\lambda g_{\tau}(\lambda)}+\frac{\lambda g_{\tau}(\lambda)}{2}.$
@@
For a fixed $t$, if $\lambda$ is such that $\lambda g_{t}(\lambda)=\sqrt{2D_{t}(\delta)}$ (which is possible in the sub-Gaussian case since the map $x\mapsto xg_{\psi_{N},t}(x)$ is surjective on $\mathbb{R}_{\geq 0}$), we obtain $\|S_{t}\|_{(V_{t}+U_{0})^{-1}}^{2}\leq D_{t}(\delta)$. This matches Theorem 3.1 up to an additive factor of 2. However, this only demonstrates that there exists a $\lambda$ which makes the widths roughly equal. We cannot in fact choose $\lambda$ such that $\lambda g_{t}(\lambda)=\sqrt{2D_{t}(\delta)}$ since $\lambda$ cannot be data-dependent. Moreover, such a choice of $\lambda$ would only ensure the widths are comparable at a particular fixed time $t$.
Remark 4.3.
That $\lambda$ is bounded away from 0 in Theorem 4.1 may seem like a cause for concern. In many non-self-normalized concentration inequalities (e.g., traditional Hoeffding, Bernstein, Bennett bounds) one typically chooses $\lambda$ as a decreasing function of $n$. However, that is not the case with the self-normalized inequalities studied here, whose widths are growing with time. For this reason, the optimal value of $\lambda$ at a particular time $t=n$ is an increasing function of $n$ (this is illustrated concretely by the preceding discussion, and by the proof of Theorem B.1).
@@
If $\lambda$ is not optimized, how does Theorem 4.1 with $\psi=\psi_{N}$ compare to Theorem 3.1? The right hand panel of Figure 1 plots the comparison for various values of $\lambda$. As one would expect, Theorem 4.1 is looser since it handles a wider class of distributions and is not explicitly optimized for sub-Gaussian processes. That said, Theorem 4.1 loses surprisingly little over Theorem 3.1, even for fixed values of $\lambda$. However, $\lambda$ must be chosen correctly, otherwise the bound can become extremely poor. As we can see, one faces a tradeoff when choosing $\lambda$, with smaller values being tighter for larger $t$ (though sometimes this tradeoff is minimal compared to the size of the bound; see Figure 2 for instance). The early staggered jumps that we see in the sub-Gaussian boundary in Figure 1 are a result of small values of $g_{\psi,t}$ early on, before $\alpha_{t}$ shrinks sufficiently.
Figure 2 compares Theorem 4.1 to the bound presented by Whitehouse et al. [62]. We fix the growth rate of the condition number, $\kappa(V_{t})$, and vary that of $\det(V_{t})$. As expected, for most values of $\lambda$, our bound outperforms the condition number-based bound of Whitehouse et al. [63] when the determinant grows slowly but is uniformly weaker when the determinant grows quickly.
Intriguingly, the bound of Whitehouse et al. [63] seems to always dominate our bound at small sample sizes (roughly <200<200). The is due to the factor of $g_{\psi,t}(\lambda)$ in our bounds, which is small at low sample sizes (or, more specifically, until $\gamma_{\min}(V_{t}+U_{0})$ grows sufficiently large). This highlights a drawback of our bounds: if $\gamma_{\min}(V_{t}+U_{0})$ never grows, or grows extremely slowly, then $g_{\psi,t}(\lambda)$ can remain extremely small, blowing up the bound and making it vacuous.
Next we explore how to remove $\lambda$ as a hyperparameter in Theorem 4.1.
Figure 2: A comparison of Theorem 4.1 with the bound of Whitehouse et al. [63], which we recall is based on the condition number of $V_{\tau}$. We control the growth of the determinant with rank-$k$ updates: as $k$ grows, so does $\det V_{\tau}$. The condition number has the same growth in each case. As the sample size grows, Theorem 4.1 outperforms the condition number-based bound for smaller values of $\det V_{\tau}$, though the performance depends substantially on the choice of $\lambda$. It loses ground as the determinant grows relative to $d\log\kappa(V_{\tau})$. We use $d=20,c=1$ and $U_{0}=I_{d}$. Full simulation details can be found in Appendix C.